To make a solution, 38.7 g of a nonelectrolyte is dissolved into 218 g of water. The freezing point of the solution is measured to be -5.53°C. Calculate the molar mass of the solute.

1
Expert's answer
2022-03-12T02:00:42-0500

∆Tf = iKfm

For water, Kf=1.86oc/kg/mol

For a non-electrolyte, i=1

m=molality of solution

m=(38.7×1000)÷(218×molar mass)

= 177.5229÷molar mass


5.53=1×1.86×177.5229÷molar mass

Molar mass= 1.86×177.5229÷5.53

=59.71